40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 805     Entry time: Wed Aug 6 19:01:15 2008
Author: Alberto 
Type: Update 
Category: General 
Subject: ITMX and SRM OSEM post-earthquake diagnostic 
Koji, Yoichi, Alberto

Today we reset the OSEMs on ITMX and SRM in order to be centered when the mirrors are aligned to the IFO beam. Since the PRM is still out of order, we used the beam from NPRO laser of the absolute length measurement experiment as it is injected through the AS port.
That’s how we did it:

1) We aligned the SRM so that the reflected beam from the NPRO was at the camera after at the AS port.

2) We traded off the alignment of SRM in order for the reflected beam at the camera to have a nice shape, avoiding any clipping from the optics, and for the optical lever to be not too far from zero. The final alignment for SRM, as read on the sliders on the MDM screen, is: Pitch=1.1650, Yaw=1.4674.

3) We aligned ITMX checking out by an IR card that the incoming and the reflected main beam in between ITMX and the BS matched. The alignment of the two beams was improved checking the matching after the SRM. The final alignment for ITMX, as read on the sliders on the MDM screen, is: Pitch=-1.2937, Yaw=-0.9890.

4) After the alignment of SRM and ITMX these were the voltages at the OSEMs:

SRM
UL=0.957
UR=1.254
LR=0.768
LL=0.620
Side=0.958

ITMX
UL=1.144
UR=1.360
LR=0.591
LL=0.325
Side=-----

5) Finally we centered the OSEMs on both mirrors and we read these voltages:

SRM
UL=0.939
UR=0.994
LR=0.782
LL=0.938
Side=0.953

ITMX
UL=0.918
UR=0.891
LR=0.887
LL=0.875
Side=0.883
ELOG V3.1.3-